# Secure numerical simulations using fully homomorphic encryption

### v2

@@ 3.3 Encryption and decryption
As shown in Fig. 4(a), the error of a single encryption/decryption operation is of $\mathcal{O}(10^{-13})$, and remains approximately there even after more then ten subsequent encryption/decryption pairs. Since the error is approximately in the range of machine precision (for the parameters summarized in Table 1), and usually only a single encryption and decryption are performed, the error is negligible for all practical purposes.
@@
Please note that due to the nature of how we measure the accuracy of other operations, the error incurred by one encryption/decryption operation will be included in the error analysis of all other FHE operations in the subsequent sections. Moreover, since it is impossible to measure the error of encoding and decoding operations independently, they are also implicitly included. A more detailed discussion of encryption/decryption errors can be found in [44].
@@ 3.4 Addition
When adding ciphertexts, it is instructive to distinguish between ciphertexts with correlated and uncorrelated errors. To achieve uncorrelated ciphertext errors, one needs to encrypt the second ciphertext summand independently for each operation (or rerandomize the ciphertext by adding an encryption of zero), whereas for correlated ciphertext errors, the same ciphertext is added to itself. As shown in Fig. 5(a), the difference between these two cases is significant. Although the errors initially have the same magnitude, correlated errors increase linearly with the number of additions, as the same errors are repeatedly added. In contrast, uncorrelated errors grow at a rate proportional to the square root of the number of additions, matching the expected behavior based on the Central Limit Theorem.
@@ 3.5 Multiplication
Next, we analyze the accuracy and performance of multiplication operations. As shown in Fig. 6(a), the behavior of the error from multiplying two ciphertexts depends on whether the ciphertexts have correlated or uncorrelated errors. Similar to the findings for ciphertext addition, multiplying ciphertexts with correlated errors leads to a linear increase in overall error, while multiplying ciphertexts with uncorrelated errors results in an error growth proportional to the square root of the number of operations. In both cases, we adjusted the levels of the ciphertexts used for multiplication to maintain consistency across the ciphertexts. In practice, the total error from multiplication in an FHE algorithm will be between the correlated and uncorrelated error curves.
@@
As before, the runtime grows with increased multiplicative depth $l_{\text{max}}$. Note that for the performance measurements, we evaluated the runtime of the second multiplication of the ciphertext with another factor, since in OpenFHE with the FLEXIBLEAUTO rescaling technique, the rescaling step is not applied to the first multiplication.
@@ 3.7 Bootstrapping
From Fig. 8(a), it becomes clear that bootstrapping has a much higher approximation error than all other CKKS operations, which is consistent with previous findings, e.g., in [41]. The $L^{\infty}$ error at $\mathcal{O}(10^{-6})$ for standard bootstrapping and $\mathcal{O}(10^{-10})$ for iterative bootstrapping is orders of magnitude greater than in all previous experiments (note that adding further bootstrapping iterations is expected to make this error comparable to the error after leveled CKKS operations). When performing multiple bootstrapping procedures in a row, the error seems to remain approximately constant. This may be specific to a relatively small value of the batch size $L=64$ and data points chosen for our experiments (note that in [41] a gradual decline in precision is observed, reaching a plateau for a higher number of bootstrapping operations). Fig. 8(b) further reveals that bootstrapping is also the most time-consuming of all CKKS operations, with a runtime that again grows linearly with the maximum multiplicative depth $l_{\text{max}}$. The runtime for the iterative bootstrapping is approximately twice as large as for the standard bootstrapping. Thus in practice, the increased accuracy of the iterative bootstrapping technique has to be balanced with the additional computational cost. Furthermore, we have to emphasize that the errors and runtimes reported here are highly dependent on the specific CKKS setup and the actual user data, and should thus be only interpreted as indicators and not taken for their absolute values. Unless noted otherwise, in the remainder of the paper we will use the standard bootstrapping method.
@@
At this point it should be mentioned that the bootstrapping procedure (as implemented in OpenFHE) requires one remaining multiplicative level per bootstrapping iteration to get started and also consumes multiple levels for the bootstrapping itself. Therefore, after bootstrapping the ciphertext level is restored to $l=l_{\text{refresh}}<l_{\text{max}}$, and the number of levels for continuous operation is $l_{\text{usable}}=l_{\text{refresh}}-1$ for standard and $l_{\text{usable}}=l_{\text{refresh}}-2$ for iterative bootstrapping. Fig. 9 illustrates these values, which are important to consider when designing an FHE application.
Remark 3.1.
Please note that in the OpenFHE library, the value returned by the GetLevel() function does not represent the ciphertext level as it is commonly used in literature and in this paper. Instead, it starts at zero for a fresh ciphertext and increases by one with each multiplication. Also, when using the OpenFHE library with the FLEXIBLEAUTO scaling technique as we are here (see Table 1 above), the rescaling step after a ciphertext multiplication is not performed immediately but only just before the next multiplication. Therefore, the ciphertext level reported by GetLevel() function may be off by one. These implementation-specific peculiarities must be taken into account when implementing algorithms with OpenFHE.
@@ 3.8 Overview of accuracy and performance of OpenFHE operations
As noted at the beginning of the accuracy and performance investigation, the exact values of the errors and execution times depend on many factors such as the chosen CKKS parameters, the actual operations performed, the data itself, and on the used compute system. Consequently, all absolute values presented here should be regarded as indicative estimates. We summarize our findings in Table 2, where we analyze the errors and execution times of different FHE operations relative to ciphertext-ciphertext addition. From the table, it is again evident that the most inaccurate and time-consuming FHE operation is bootstrapping. Hence, it is important to use bootstrapping as rarely as possible. On the other hand, the previous results have also shown a linear dependence of the runtime on the maximum multiplicative depth $l_{\text{max}}$ for all FHE operations. Although increasing $l_{\text{max}}$ and the usable levels after bootstrapping $l_{\text{usable}}$ reduces the number of required bootstrapping operations, the runtime of all operations (including bootstrapping itself) increases linearly. Therefore, $l_{\text{max}}$ should be chosen carefully to balance error and runtime, and its optimal value depends on the specific application.
@@
Encrypt plaintext $<1$ $\approx 20$
Decrypt ciphertext $<1$ $\approx 30$
Ciphertext + ciphertext $=1$ $=1$
Ciphertext + plaintext $<1$ $\approx 1$
Ciphertext + scalar $<1$ $<1$
@@
Ciphertext * plaintext $<1$ $\approx 10$
Ciphertext * scalar $<1$ $\approx 6$
Rotate ciphertext $<1$ $\approx 50$
@@ 4 Setting up secure numerical simulations
In this section, we will begin by introducing the linear advection equation in one and two spatial dimensions, which will serve as the prototypical PDEs for our secure numerical simulations. We will then present two finite difference schemes that we will use to discretize these equations. Next, we will rewrite these schemes in terms of the FHE primitives provided by the CKKS scheme. Finally, we will provide the full algorithm for secure numerical simulations using FHE.
@@ 4.1 Linear scalar advection equations
The linear scalar advection equation describes the transport of a scalar field $u$ with a constant speed. In one dimension, it is given by
[…]
where $u=u(t,x)$ is the scalar field, $t$ is time, $x$ is the spatial coordinate, and $a_{x}$ is the advection speed in the $x$-direction. In two spatial dimensions, the linear scalar advection equation reads
[…]
where we now have $u=u(t,x,y)$, $y$ as the second spatial coordinate, and $a_{y}$ as the advection speed in the $y$-direction.
For the simulations, we will use the computational domain $\Omega=[0,1]$ in 1D and $\Omega=[0,1]\times[0,1]$ in 2D. To simplify the FHE implementation, we will consider periodic boundary conditions in all spatial directions. Furthermore, we will restrict our case to positive advection speeds, i.e., $a_{x},a_{y}>0$. A suitable initial condition will be given later for each simulation.
@@ 4.3 Reformulating finite difference schemes for FHE
All FHE operations in the CKKS scheme are vector-based, thus we need to rewrite the finite difference schemes in Eqns. (10)–(13) in a vector-wise manner. Furthermore, we need to limit ourselves to the available operations of element-wise addition, multiplication, and rotation. We will begin with the 1D equations and then proceed to the 2D equations. In the latter case, we will also need to develop a strategy to handle matrix operations with FHE.
In the following, we will again use bold symbols to denote vectors and matrices, e.g., $\bm{u}=(u_{1},u_{2},...,u_{M})^{\intercal}$. Vector/matrix addition and subtraction are written using the usual symbols $+$ and --, respectively. Element-wise vector/matrix multiplication is indicated by the Hadamard product symbol $\odot$. In the spirit of the SecureArithmetic.jl library, we will not distinguish in our notation between unencrypted and encrypted data, since all algorithms work transparently with either.
@@
Depending on the context, the “$\operatorname{circshift}$” operator may represent either the rotation of unencrypted or encrypted vectors and matrices. When we specifically refer to the CKKS rotation operation, we will use the term “$\operatorname{rotate}$”. For our goal to rewrite the finite difference schemes in terms of operations supported by FHE, we thus have addition/subtraction, multiplication, and $\operatorname{circshift}$ at our disposal.
@@ 4.3.1 Rewriting the finite difference methods in 1D
Recall from Sec. 2.1.1 that in the CKKS scheme, vectors of real numbers are encoded into plaintexts with a certain batch size/capacity, which is at most half the ring dimension $N_{R}$. If the batch size coincides with the length of the solution vector $\bm{u}$, we can directly use the CKKS rotation operation for our $\operatorname{circshift}$ operation. In general, however, we need a way to handle the rotation of vectors where the length of the user-provided vector is smaller than the capacity. This is especially relevant for practical applications, since the batch size is always a power of two.
Therefore, we need to augment the $\operatorname{circshift}$ operation to support the case where the length of the actual data vector is less than the capacity of the ciphertext. We present an algorithm for this in Alg. 1. The algorithm performs two rotations of the input ciphertext $\bm{x}$, one backward and one forward, and then uses multiplication with appropriate masking vectors to combine the elements from both rotations. In the algorithm, we designate the CKKS rotation operation as “$\operatorname{rotate}$”. Since the OpenFHE library uses a different sign convention for the rotation index than we use for the $\operatorname{circshift}$ operation (which follows the sign convention of the corresponding function in the Julia base library, Base.circshift), we will need to adjust the sign of the shift in the $\operatorname{rotate}$ operation. Note that for the input vectors in Alg. 1, we consider the length to be equal to the length of the actual user data, while the capacity is equal to the batch size.
@@
2:
3:$\bm{u}\leftarrow\operatorname{rotate}(\bm{x},-k)$
[…]
6:$f_{1}\leftarrow 1$
7:$l_{1}\leftarrow\text{length}(x)+k$
8:$f_{2}\leftarrow 1+\text{length}(x)+k$
9:$l_{2}\leftarrow\text{length}(x)$Else
10:$f_{1}\leftarrow 1+k$
11:$l_{1}\leftarrow\text{length}(x)$
$f_{2}\leftarrow 1$
13:$l_{2}\leftarrow k$EndIf
14:
15:$\bm{m}_{1}\leftarrow(0,\dots,0)^{\intercal}$triangleright Create masking vectors as plaintexts
[…]
17:$\bm{m}_{2}\leftarrow(0,\dots,0)^{\intercal}$
[…]
19:
20:$\bar{\bm{u}}\leftarrow\bm{u}\odot\bm{m}_{1}$$\triangleright$ Apply masks by multiplying ciphertexts with plaintexts
21:$\bar{\bm{v}}\leftarrow\bm{v}\odot\bm{m}_{2}$
22:
[…]
With this implementation for the $\operatorname{circshift}$ operation, we now support all operations to compute the solution at the new time step in Eqns. (16) and (17). Note that $\operatorname{circshift}$ consumes one multiplicative level if the length of the ciphertext is less than its capacity, and zero levels otherwise. Therefore, a single evaluation of Eqn. (16) or Eqn. (17) uses two multiplicative levels if the length of the ciphertext $\bm{u}$ is less than the capacity, and one level otherwise. Alg. 1 is also the basis for the $\operatorname{circshift}$ implementation in the SecureArithmetic.jl package.
@@ 4.3.2 Matrix arithmetic with FHE ciphertexts
We therefore need to extend Alg. 1 to two dimensions. One possible approach is given in Alg. 2. Similar to the 1D algorithm, it works through a series of rotations and appropriate masking vectors, which are then additively combined for the final result. With the “size” function we obtain the number of rows and columns of the matrix. Note that the two-dimensional $\operatorname{circshift}$ operation consumes between zero and two multiplicative levels, depending on whether a mask is required and if the underlying ciphertext holding the matrix data has a length that is equal or less than its capacity. An overview of the required number of multiplications is given in Table 3. Note that Alg. 2 is also the basis for the $\operatorname{circshift}$ implementation for ciphertexts representing matrix data in the SecureArithmetic.jl package.
@@
[…]
2:return $\operatorname{circshift}(\bm{x},ln)$EndIf
[…]
4:$f_{1}\leftarrow 1$
5:$l_{1}\leftarrow n-k$
6:$f_{2}\leftarrow 1+n-k$
7:$l_{2}\leftarrow n$Else
8:$f_{1}\leftarrow 1-k$
9:$l_{1}\leftarrow n$
10:$f_{2}\leftarrow 1$
11:$l_{2}\leftarrow-k$EndIf
12:
13:$\bm{m}_{\text{tmp},1}\leftarrow(0,0,\dots,0)^{\intercal}\in\mathbb{R}^{n}$triangleright Create masking vectors for one column as plaintexts
[…]
15:$\bm{m}_{\text{tmp},2}\leftarrow(0,0,\dots,0)^{\intercal}\in\mathbb{R}^{n}$
[…]
17:$\bm{m}_{1}\leftarrow$repeat$(\bm{m}_{\text{tmp},1},m)\in\mathbb{R}^{nm}$triangleright Repeat the mask for each column
18:$\bm{m}_{2}\leftarrow$repeat$(\bm{m}_{\text{tmp},2},m)\in\mathbb{R}^{nm}$
19:
20:$\bar{\bm{u}}\leftarrow\bm{x}\odot\bm{m}_{1}$$\triangleright$ Apply masks by multiplying ciphertexts with plaintexts
21:$\bar{\bm{v}}\leftarrow\bm{x}\odot\bm{m}_{2}$
22:
23:$s_{1}\leftarrow ln+k$triangleright Compute one-dimensional shift indices
[…]
26:return $\operatorname{rotate}(\bar{\bm{u}},-s_{1})+\operatorname{rotate}(\bar{\bm{v}},-s_{2})$Else
[…]

### v3

@@ 3.3 Encryption and decryption
As shown in Fig. 6(a), the error of a single encryption/decryption operation is of $\mathcal{O}(10^{-13})$, and remains at that level even after more than ten subsequent encryption/decryption pairs. Since the error is approximately in the range of machine precision (for the parameters summarized in Table 1), and usually only a single encryption and decryption are performed, the error is negligible for all practical purposes.
@@
Due to the nature of how we measure the accuracy of other operations, the error incurred by one encryption/decryption operation will be included in the error analysis of all other FHE operations in the subsequent sections. Moreover, since it is impossible to measure the error of encoding and decoding operations independently, they are also implicitly included. A more detailed discussion of encryption/decryption errors can be found in [45].
@@ 3.4 Addition
When adding ciphertexts, it is instructive to distinguish between ciphertexts with correlated and uncorrelated errors. Two ciphertexts have correlated errors if they share the same noise or if their noise terms are not statistically independent. This typically happens when the same ciphertext is used more than once in an algorithm, e.g., when adding a ciphertext to itself multiple times. To achieve uncorrelated ciphertext errors, one needs to encrypt the second ciphertext summand independently for each operation (or rerandomize the ciphertext by adding an encryption of zero). As shown in Fig. 7(a), the difference between these two cases is significant. Although the errors initially have the same magnitude, correlated errors increase linearly with the number of additions, as the same errors are repeatedly added. In contrast, uncorrelated errors grow at a rate proportional to the square root of the number of additions, matching the expected behavior based on the Central Limit Theorem.
@@ 3.5 Multiplication
Next, we analyze the accuracy and performance of multiplication operations. As shown in Fig. 8(a), the behavior of the error from multiplying two ciphertexts depends on whether the ciphertexts have correlated or uncorrelated errors. Similar to the findings for ciphertext addition, multiplying ciphertexts with correlated errors leads to a linear increase in the overall error, while multiplying ciphertexts with uncorrelated errors results in an error growth proportional to the square root of the number of operations. In both cases, we adjusted the levels of the ciphertexts used for multiplication to maintain consistency across the ciphertexts. In practice, the total error from multiplication in an FHE algorithm will be between the correlated and uncorrelated error curves.
@@
As before, the runtime grows with increased multiplicative depth $l_{\text{max}}$. For the performance measurements, we evaluated the runtime of the second multiplication of the ciphertext with another factor, since in OpenFHE with the FLEXIBLEAUTO rescaling technique, the rescaling step is not applied to the first multiplication.
@@ 3.7 Bootstrapping
From Fig. 10(a), it becomes clear that bootstrapping has a much higher approximation error than all other CKKS operations, which is consistent with previous findings, e.g., in [42]. The $L^{\infty}$ error at $\mathcal{O}(10^{-6})$ for standard bootstrapping and $\mathcal{O}(10^{-9})$ for iterative bootstrapping is orders of magnitude greater than in all previous experiments (adding further bootstrapping iterations is expected to make this error comparable to the error after leveled CKKS operations). When performing multiple bootstrapping procedures in a row, a gradual decline in precision is observed, with the rate of precision loss becoming progressively smaller. This is consistent with [42], where the error eventually reaches quasi-steady state conditions after many bootstrapping invocations. The primary cause for this behavior is the approximation of the modular reduction—the function evaluated as part of CKKS bootstrapping—with a sine wave. This approximation is only accurate near zero within a small, periodically repeating interval, and its error grows as the message value moves farther from zero [42]. Each application of the sine wave approximation introduces additional error, but the precision loss decreases with each step, as each bootstrapping operates on already-degraded input (rather than the more accurate value fed into the first bootstrapping). In practice, this means that CKKS can be used for very deep computations with many hundreds of bootstrapping invocations, with a small cumulative loss in precision compared to the first bootstrapping. Moreover, CKKS also provides tools to further reduce the noise using Hermite interpolations [53], thereby enabling practically unlimited deep computations.
Fig. 10(b) further reveals that bootstrapping is also the most time-consuming of all CKKS operations, with a runtime that again grows linearly with the maximum multiplicative depth $l_{\text{max}}$. The runtime for the iterative bootstrapping is approximately twice as large as for the standard bootstrapping. Thus in practice, the increased accuracy of the iterative bootstrapping technique has to be balanced with the additional computational cost. Furthermore, we have to emphasize that the errors and runtimes reported here are highly dependent on the specific CKKS setup and the actual user data, and should thus be only interpreted as indicators and not taken for their absolute values. Unless noted otherwise, in the remainder of the paper we use the standard bootstrapping method.
@@
At this point it should be mentioned that the bootstrapping procedure (as implemented in OpenFHE) requires one remaining multiplicative level per bootstrapping iteration to get started and also consumes multiple levels for the bootstrapping itself. Therefore, after bootstrapping the ciphertext level is restored to $l=l_{\text{refresh}}<l_{\text{max}}$, and the number of levels for continuous operation is $l_{\text{usable}}=l_{\text{refresh}}-1$ for standard and $l_{\text{usable}}=l_{\text{refresh}}-2$ for iterative bootstrapping. Fig. 11 illustrates these values, which are important to consider when designing an FHE application. A more detailed discussion of the slightly different behavior of the corresponding OpenFHE function GetLevel() is provided in A.
@@ 3.8 Overview of accuracy and performance of OpenFHE operations
As noted at the beginning of the accuracy and performance investigation, the exact values of the errors and execution times depend on many factors such as the chosen CKKS parameters, the actual operations performed, the data itself, and on the used compute system. Consequently, all absolute values presented here should be regarded as indicative estimates. We summarize our findings in Table 2, where we analyze the errors and execution times of different FHE operations relative to ciphertext-ciphertext addition. From the table, it is again evident that the most inaccurate and time-consuming FHE operation is bootstrapping. Hence, it is important to use bootstrapping as rarely as possible. At the same time, the previous results also show a linear dependence of the runtime on the maximum multiplicative depth $l_{\text{max}}$ for all FHE operations. Although increasing $l_{\text{max}}$ and the usable levels after bootstrapping $l_{\text{usable}}$ reduces the number of required bootstrapping operations, the runtime of all operations (including bootstrapping itself) increases linearly. Therefore, $l_{\text{max}}$ should be chosen carefully to balance error and runtime, and its optimal value depends on the specific application.
@@
Encrypt plaintext <1<1 $\approx 20$
Decrypt ciphertext <1<1 $\approx 30$
Ciphertext + ciphertext =1=1 =1=1
Ciphertext + plaintext <1<1 $\approx 1$
Ciphertext + scalar <1<1 <1<1
@@
Ciphertext * plaintext <1<1 $\approx 10$
Ciphertext * scalar <1<1 $\approx 6$
Rotate ciphertext <1<1 $\approx 50$
@@ 4 Setting up secure numerical simulations
In this section, we begin by introducing the linear advection equation in one and two spatial dimensions, which will serve as the prototypical PDEs for our secure numerical simulations. We then present two finite difference schemes which we use to discretize these equations. Next, we rewrite these schemes in terms of the FHE primitives provided by the CKKS scheme. Finally, we provide the full algorithm for secure numerical simulations using FHE.
@@ 4.1 Linear scalar advection equations
The linear scalar advection equation describes the transport of a scalar field $u$ at constant speed. In one dimension, it is given by
$\frac{\partial u}{\partial t}+a_{x}\frac{\partial u}{\partial x}=0,$ (8)
with $u=u(t,x)$, where $t$ is time, $x$ is the spatial coordinate, and $a_{x}>0$ is the advection speed. In two dimensions, the equation extends to
$\frac{\partial u}{\partial t}+a_{x}\frac{\partial u}{\partial x}+a_{y}\frac{\partial u}{\partial y}=0,$ (9)
with $u=u(t,x,y)$, and positive speeds $a_{x},a_{y}>0$.
Simulations are conducted on the domains $\Omega=[0,1]$ in 1D and $\Omega=[0,1]\times[0,1]$ in 2D, using periodic boundary conditions. Initial conditions are specified in the respective examples.
@@ 4.3 Reformulating finite difference schemes for FHE
All FHE operations in the CKKS scheme are vector-based, thus we need to rewrite the finite difference schemes in Eqns. (10)–(13) in a vector-wise manner. Furthermore, we need to limit ourselves to the available operations of element-wise addition, multiplication, and rotation. We begin with the 1D equations and then proceed to the 2D equations. In the latter case, we also need to develop a strategy to handle matrix operations with FHE.
In the following, vector/matrix addition and subtraction are written using the usual symbols ++ and -, respectively. Element-wise vector/matrix multiplication is indicated by the Hadamard product symbol $\odot$. In the spirit of the SecureArithmetic.jl library, we do not distinguish in our notation between unencrypted and encrypted data, since all algorithms work transparently with either.
@@
Depending on the context, the $\operatorname{circshift}$” operator may represent either the rotation of unencrypted or encrypted vectors and matrices. When we specifically refer to the CKKS rotation operation, we use the term $\operatorname{rotate}$”. For our goal to rewrite the finite difference schemes in terms of operations supported by FHE, we thus have addition/subtraction, multiplication, and $\operatorname{circshift}$ at our disposal.
@@ 4.3.1 Rewriting the finite difference methods in 1D
As described in Sec. 2.1.1, the CKKS scheme encodes vectors of real numbers into plaintexts with a certain batch size/capacity, which is at most half the ring dimension $N_{R}$. If the batch size coincides with the length of the solution vector $\bm{u}$, we can directly use the CKKS rotation operation for our $\operatorname{circshift}$ operation. In general, however, we need a way to handle the rotation of vectors where the length of the user-provided vector is smaller than the capacity. This is especially relevant for practical applications, since the batch size is always a power of two.
Therefore, we need to augment the $\operatorname{circshift}$ operation to support the case where the length of the actual data vector is less than the capacity of the ciphertext. We present an algorithm for this in Alg. 1. The algorithm performs two rotations of the input ciphertext $\bm{x}$, one backward and one forward, and then uses multiplication with appropriate masking vectors to combine the elements from both rotations. In the algorithm, we designate the CKKS rotation operation as $\operatorname{rotate}$”. Since the OpenFHE library uses a different sign convention for the rotation index than we use for the $\operatorname{circshift}$ operation (which follows the sign convention of the corresponding function in the Julia base library, Base.circshift), we need to adjust the sign of the shift in the $\operatorname{rotate}$ operation. For the input vectors in Alg. 1, we consider the length to be equal to the length of the actual user data, while the capacity is equal to the batch size.
@@
4:    return $\operatorname{rotate}(\bm{x},-k)$ $\triangleright$ Use CKKS rotation if data length matches ciphertext capacity
5:   end if 6:
7:   $\bm{u}\leftarrow\operatorname{rotate}(\bm{x},-k)$
8:   $\bm{v}\leftarrow\operatorname{rotate}(\bm{u},(k>0$ ? $(\bm{x}) : -length$$(\bm{x}))-k)$ 9:
10:   if $k<0$ then $\triangleright$ Determine indices for masking vectors
11:    $f_{1}\leftarrow 1$
12:    $l_{1}\leftarrow\text{length}(x)+k$
13:    $f_{2}\leftarrow 1+\text{length}(x)+k$
14:    $l_{2}\leftarrow\text{length}(x)$
15:   else
16:    $f_{1}\leftarrow 1+k$
17:    $l_{1}\leftarrow\text{length}(x)$
18:    $f_{2}\leftarrow 1$
19:    $l_{2}\leftarrow k$
20:   end if 21:
22:   $\bm{m}_{1}\leftarrow(0,\dots,0)^{\intercal}$ $\triangleright$ Create masking vectors as plaintexts
23:   $\bm{m}_{1}[f_{1}:l_{1}]\leftarrow 1$
24:   $\bm{m}_{2}\leftarrow(0,\dots,0)^{\intercal}$
25:   $\bm{m}_{2}[f_{2}:l_{2}]\leftarrow 1$ 26:
27:   $\bar{\bm{u}}\leftarrow\bm{u}\odot\bm{m}_{1}$ $\triangleright$ Apply masks by multiplying ciphertexts with plaintexts
28:   $\bar{\bm{v}}\leftarrow\bm{v}\odot\bm{m}_{2}$ 29:
30:   return $\bar{\bm{u}}+\bar{\bm{v}}$ $\triangleright$ Combine masked ciphertexts for final result
31: end function
With this implementation for the $\operatorname{circshift}$ operation, we now support all operations to compute the solution at the new time step in Eqns. (16) and (17). The $\operatorname{circshift}$ operation consumes one multiplicative level if the length of the ciphertext is less than its capacity, and zero levels otherwise. Therefore, a single evaluation of Eqn. (16) or Eqn. (17) uses two multiplicative levels if the length of the ciphertext $\bm{u}$ is less than the capacity, and one level otherwise. Alg. 1 is also the basis for the $\operatorname{circshift}$ implementation in the SecureArithmetic.jl package.
@@ 4.3.2 Matrix arithmetic with FHE ciphertexts
We therefore need to extend Alg. 1 to two dimensions. One possible approach is given in Alg. 2. Similar to the 1D algorithm, it works through a series of rotations and appropriate masking vectors, which are then additively combined for the final result. With the “size” function we obtain the number of rows and columns of the matrix. The two-dimensional $\operatorname{circshift}$ operation consumes between zero and two multiplicative levels, depending on whether a mask is required and if the underlying ciphertext holding the matrix data has a length that is equal or less than its capacity. An overview of the required number of multiplications is given in Table 3. Alg. 2 is also the basis for the $\operatorname{circshift}$ implementation for ciphertexts representing matrix data in the SecureArithmetic.jl package.
@@
2: function circshift($\bm{x}$, $k$, $l$)
3:   $n,m=\text{size}(\bm{x})$ $\triangleright$ Get matrix dimensions $n\times m$
4:   if $k==0$ then $\triangleright$ If no row shift, directly use vectorial $\operatorname{circshift}$ from Alg. 1
5:    return $\operatorname{circshift}(\bm{x},ln)$
6:   end if 7:
8:   if $k>0$ then $\triangleright$ Determine indices for masking vectors
9:    $f_{1}\leftarrow 1$
10:    $l_{1}\leftarrow n-k$
11:    $f_{2}\leftarrow 1+n-k$
12:    $l_{2}\leftarrow n$
13:   else
14:    $f_{1}\leftarrow 1-k$
15:    $l_{1}\leftarrow n$
16:    $f_{2}\leftarrow 1$
17:    $l_{2}\leftarrow-k$
18:   end if 19:
20:   $\bm{m}_{\text{tmp},1}\leftarrow(0,0,\dots,0)^{\intercal}\in\mathbb{R}^{n}$ $\triangleright$ Create masking vectors for one column as plaintexts
21:   $\bm{m}_{\text{tmp},1}[f_{1}:l_{1}]\leftarrow 1$
22:   $\bm{m}_{\text{tmp},2}\leftarrow(0,0,\dots,0)^{\intercal}\in\mathbb{R}^{n}$
23:   $\bm{m}_{\text{tmp},2}[f_{2}:l_{2}]\leftarrow 1$
24:   $\bm{m}_{1}\leftarrowrepeat$$(\bm{m}_{\text{tmp},1},m)\in\mathbb{R}^{nm}$ $\triangleright$ Repeat the mask for each column
25:   $\bm{m}_{2}\leftarrowrepeat$$(\bm{m}_{\text{tmp},2},m)\in\mathbb{R}^{nm}$ 26:
27:   $\bar{\bm{u}}\leftarrow\bm{x}\odot\bm{m}_{1}$ $\triangleright$ Apply masks by multiplying ciphertexts with plaintexts
28:   $\bar{\bm{v}}\leftarrow\bm{x}\odot\bm{m}_{2}$ 29:
30:   $s_{1}\leftarrow ln+k$ $\triangleright$ Compute one-dimensional shift indices
31:   $s_{2}\leftarrow ln+k+(k<0$ ? $n$ : n)-n) 32:
33:   if $l==0$ then $\triangleright$ Without column shift, save multiplications by using $\operatorname{rotate}$
34:    return $\operatorname{rotate}(\bar{\bm{u}},-s_{1})+\operatorname{rotate}(\bar{\bm{v}},-s_{2})$
35:   else
36:    return $\operatorname{circshift}(\bar{\bm{u}},s_{1})+\operatorname{circshift}(\bar{\bm{v}},s_{2})$
37:   end if
38: end function